# LeapFrog: The Rowhammer Instruction Skip Attack

### v1

@@ 8 Experiments
TLS 29KB 5328007 1 315 2493
2 2240 14413
3 21841 67421
@@ 3 N/A 70475
sudo 227KB 148177 1 1100 8655
2 6910 54203
3 30908 221181
Table 1: Number of gadget candidates found by MFS in for fault models with different Hamming distances. We ran OpenSSL with aria-128-cbc cipher.
8.1 OpenSSL Encryption Bypass
We analyze openssl command line tool that uses OpenSSL v1.1.1w for block cipher and stream cipher implementations. For each cipher, we give a simple plaintext that contains the helloworld string and run encryption without salt with a simple passphrase. Our aim is to find LeapFrog gadgets in the binary that can be exploited for bypassing encryption steps in the ciphers, revealing the plaintext.
First, we scan the binary using MFS as described in §7. Since we do not aim for any authentication bypass in this scenario, and the execution traces are deterministic for fixed inputs, step 2 is not applicable. Instead, in step 3, we compare the return addresses in a single trace against all the instruction addresses in the same trace to look for targets with $d_{HD}=1$.
Fig. 8 illustrates one of the LeapFrog gadgets found in the openssl command line tool. When we corrupt a single bit in 0x55555559c4c5, the return address of opt_cipher function, to make it 0x55555559c0d5, the function returns to the corrupted return address, skipping three instructions in between. Similarly, another single-bit corruption to (0x55555559c0c5) causes the function to return to an earlier point in the program. We verified that both of these bit flips cause the binary to skip the whole encryption and instead output the plaintext. Similarly, MFS detected LeapFrog gadgets that are used in 36 ciphers including block ciphers and stream ciphers. The ciphers with LeapFrog gadgets that revealed full or partial plaintext are listed in Table 2. Fig. 6 and  7 summarize the simulation results for aes-256-ctr and aria-256-ctr respectively.
@@ 8.1 OpenSSL Encryption Bypass
Table 2: 36 ciphers implemented in OpenSSL that are vulnerable to LeapFrog attack. Each given cipher reveals the plaintext fully or partially in the ciphertext due to skipped encryption steps.
[…]
8.2 Sudo Privilege Escalation
We analyze the sudo binary installed on the system for potential LeapFrog gadgets. Our objective is to bypass native password authentication implementation in sudo v1.9.9 resulting in privilege escalation.
First, we create a dummy C program that tries to bind to port 80 in the system, which is a privileged operation. As an unprivileged user, when we try to execute the program it outputs Bind failed: Permission denied error and exits. When we run the same program with root privileges, it outputs Successfully bound to port 80 and exits. The purpose of creating such a dummy program is to have a probe on the simulation that detects the attack’s success.
Then we run the program with sudo twice, first with the correct password and second with an incorrect password. This results in two different traces. We first run our detection tool with step 2 on, where we restrict ourselves to only a subset of instructions. With this mode, our tool we found 1100 unique gadget candidates. However, after running the rest of the filtering steps, none of them were found to be an actual working gadget.
Next, we turned off step 2 to increase the coverage. Our detection tool found 8655 unique candidates with $d_{HD}=1$ within the sudo binary and other dynamically linked libraries combined. After simulating these candidates, we found that 10 unique $addr_{src},addr_{dest}>$ pairs result in privilege escalation. We illustrate three of these address pairs in Fig. 9. Interestingly, $addr_{src}$ in those three gadgets are the same PC value (0x555555579970) which is the return address of the <sudo_strtoid_v2@plt> function. This shows that sudo will return to a point that allows unprivileged users to run as root with an incorrect password in case the malicious user flips one of the three identified bits during the execution of the said function. The remaining 7 gadgets we identified appear during the call instructions to <strncmp@plt> (5 times), <sudo_parse_gids_v1@plt> (1 time, also visible in Fig. 9), and <sudo_debug_enter_v1@plt> (1 time) functions. Note that stncmp function is implemented in libc library. This is a real-world example of a LeapFrog gadget caused by a combination of the instruction order in the main program binary and the PC storage in the stack during a third-party function execution. Therefore, LeapFrog mitigations should consider the role of library dependencies as well. Since a potential mitigation should be generic and independent of the program logic, the analysis of how these specific corruptions cause privilege escalation is not relevant and we leave it outside of the scope of this work.
@@ 8.2 Sudo Privilege Escalation
Figure 9: LeapFrog gadgets detected in sudo binary. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault injected, $addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
@@ 8.3 TLS Handshake
Figure 10: Probability distribution of bait page numbers.
In a full end-to-end attack example, we illustrate the potency of the attack by applying it within a client/server authentication framework, specifically using OpenSSL for signature verification. Here, we consider a scenario where the attacker shares a physical compute space with the client. The goal of the attacker is to manipulate the client’s signature verification mechanism, causing it to erroneously validate a corrupted signature as genuine. This manipulation forms part of a broader man-in-the-middle strategy, aimed at deceiving the client into believing they are securely connected to the intended server.
@@
We scan the client binary using MFS while the server is using the correct and incorrect private key. With step 2 on, we found 315 unique gadget candidates with $d_{H}=1$. When the server uses the correct key, the client binary terminates with return code 0, and when the server uses an incorrect key, the client returns with code 1. In step 5, we look for PC corruptions that cause the client to return with value 0, meaning it incorrectly authenticates the server. After simulation, we found that one of the candidates was a LeapFrog gadget that caused false authentication of the malicious server with an incorrect key.
Then, we scan the client with step 2 off. With this mode, MFS detected 2493 gadget candidates with $d_{H}=1$. After the simulation steps, we verified that 21 of those candidates were LeapFrog gadgets that cause the client to return with 0, including the one found earlier. The number of candidates for different Hamming distance values are given in Table 1.
The total time for the end-to-end attack to induce a successful misauthentication of the TLS handshake was 12 hours and 25 minutes, as seen in Table  3. This time included profiling the system for the proper flippy pages with the correct offset, meaning the actual online time was around 2 hours. The experiment found a total of 1647 unique flippy pages, and over the course of the 2 hours of online attacking, we saw 2206 attacks where the program counter was baited into the correct page we were attacking before it flipped.
@@
Table 3: Results from the end-to-end attack on code using OpenSSL client/server signature verification with LeapFrog gadget
@@ 9 Countermeasures
9.1 Rowhammer Resistant Hardware
@@
Increasing the DRAM refresh rate is a commonly cited countermeasure to prevent Rowhammer attacks. Standard DRAM refresh is 64ms, meaning that a Rowhammer attack has 64ms to flip a bit before the row refreshes. Thus a faster refresh rate will result in a shorter time window for the Rowhammer attack to be performed and should result in fewer flips. This is not an ideal solution however, because a faster refresh rate will lead to worse power usage and performance overall. Alternative methods such as probabilistic row refresh [49] and parallel row refresh [52] are not available in consumer systems.
@@
A novel countermeasure against Rowhammer attacks is the Randomized Row-Swap (RRS) method [42]. This approach fundamentally disrupts the spatial connection between aggressor and victim DRAM rows, thereby offering a robust defense against complex Rowhammer access patterns, including those not mitigated by victim-focused methods like the Half-Double attack. RRS operates by periodically swapping aggressor-rows with randomly selected rows within the DRAM memory, limiting the potential damage to any single locality. While RRS can be implemented in conjunction with any tracking mechanism, its effectiveness has been demonstrated when paired with a Misra-Gries tracker, targeting a Row Hammer Threshold of 4.8K activations, akin to state-of-the-art attacks. Notably, evaluations of RRS indicate a negligible slowdown (averaging 0.4%), while providing strong security assurances against Row Hammer bit flips even under continuous attack scenarios spanning several years.
@@
Initial beliefs held that Error Correcting Code (ECC) would serve as an effective defense against Rowhammer attacks. However, subsequent research has shown that ECC, despite its prevalence in server environments, falls short as a comprehensive solution. This inadequacy primarily arises due to ECC’s vulnerability to scenarios involving triple bit flips, a phenomenon well-documented in the literature [14]. Additionally, ECC, while standard in server-grade hardware, is typically absent in consumer-grade DRAM systems.
9.2 Adding nops To Code
@@
A mitigation against LeapFrog attack specifically would be patching the source code or binary such that it is no longer vulnerable. Given the single bit flip requirement of Rowhammer on the PC values, adding enough nops within the LeapFrog gadget to prevent instruction skips that only require a single bit flip would potentially mitigate the attack. Adding nop instructions to source code is not trivial when the compiler optimizations are enabled since the compiler may reorder the critical parts in a different way that make the patch ineffective. A mitigation tool that adds nops to binary itself may overcome the compiler effect. Yet, adding new instructions to a binary will result in a change in the address of all the following instruction which may introduce new LeapFrog gadgets. Therefore, the patched binary needs to be re-evaluated if it still has gadgets. Although, it may potentially generate a LeapFrog proof binary, we claim it is not a sound and reliable approach.
@@ 10 Conclusion
In this work, we introduced LeapFrog, a specific type of Rowhammer exploit that directly targets the control flow of programs by manipulating the Program Counter stored in the stack. This novel approach marks a significant shift in the understanding of Rowhammer threats, moving beyond traditional data integrity attacks to those that can alter program execution. Our successful demonstration of this attack in the context of an OpenSSL TLS handshake scenario highlights its practical effectiveness and potential impact on widely used security protocols.
Furthermore, we proposed a systematic approach to identify LeapFrog gadgets in real-world software. Using our MFS analysis tool, we detected multiple points in binaries of commonly used software, such as sudo and OpenSSL’s encryption tool, which results in authentication and encryption bypass, respectively, when exploited. Even though the identification of vulnerable software is relatively straightforward thanks to our detection tool, mitigation of LeapFrog is not a trivial task since it is not transparent to the developers on a source code level. Instead, dedicated Rowhammer-resistant DRAM hardware or Rowhammer-aware compiler tools will be required to prevent LeapFrog attacks.

### v2

@@ 8.1 ML Misclassification
Decision Tree 99KB 38417 1 N/A 1363
2 N/A 8667
3 N/A 32326
TABLE I: Number of gadget candidates found in decision tree algorithm with different Hamming distances.
8.2 OpenSSL Encryption Bypass
We analyze openssl command line tool that uses OpenSSL v1.1.1w for block cipher and stream cipher implementations. For each cipher, we give a simple plaintext that contains the helloworld string and run encryption without salt with a simple passphrase. We aim to find LeapFrog gadgets in the binary that can be exploited for bypassing encryption steps in the ciphers, revealing the plaintext.
First, we scan the binary using MFS as described in section 7. Since we do not aim for any authentication bypass in this scenario, and the execution traces are deterministic for fixed inputs, step 2 is not applicable. Instead, in step 3, we compare the return addresses in a single trace against all the instruction addresses in the same trace to look for targets with $d_{HD}=1$. This means that ultimately, for each return address, we test 12 different jumps, a trial for each bit in the page offset. Table II shows the total number of gadgets for each hammer distance, with a hammer distance of 1 having a total of 2700 candidates.
Target Size #Inst.exec $d_{HD}$ Candidates
2 on 2 off
@@ 3 N/A 70475
TABLE II: Number of gadget candidates found by MFS in for fault models with different Hamming distances. We ran OpenSSL with aria-128-cbc cipher.
Figure  10 illustrates one of the LeapFrog gadgets found in the openssl command line tool. When we corrupt a single bit in 0x55555559c4c5, the return address of opt_cipher function, to make it 0x55555559c0d5, the function returns to the corrupted return address, skipping three instructions in between. Similarly, another single-bit corruption to (0x55555559c0c5) causes the function to return to an earlier point in the program. We verified that both of these bit flips cause the binary to skip the whole encryption and instead output the plaintext. Similarly, MFS detected LeapFrog gadgets that are used in 36 ciphers including block ciphers and stream ciphers. The ciphers with LeapFrog gadgets that revealed full or partial plaintext are listed in Table III. Figure  8 and  9 summarize the simulation results for aes-256-ctr and aria-256-ctr respectively.
@@
TABLE III: 36 ciphers implemented in OpenSSL that are vulnerable to LeapFrog attack. Each given cipher reveals the plaintext fully or partially in the ciphertext due to skipped encryption steps.
@@
Figure 10: LeapFrog gadget in OpenSSL command line tool resulting in encryption bypass in aria-128-cbc block cipher. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault injected,$addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
8.3 Post-Quantum Cryptography Schemes
NIST announced the standards for Post-Quantum Cryptography (PQC) in FIPS 204 [48], and FIPS 205 [49]. These standards are used for digital signatures to protect against quantum attacks. We use Open Quantum Safe (liboqs version 0.11.1-dev) library [55], an open source library for PQC algorithms, to find LeapFrog gadgets on FIPS standards using MFS tool.
One of the algorithms selected by NIST for standardization is CRYSTALS-Dilithium, which serves as a digital signature scheme providing post-quantum security guarantees. Dilithium relies on the hardness of structured lattice problems, such as the Learning With Errors (LWE) problem, which is believed to be intractable for quantum computers. Another prominent algorithm is FALCON, which offers smaller key sizes and signatures by employing the NTRU lattice, making it a competitive choice for constrained environments. Our analysis of these algorithms reveals that, despite their robust design against quantum attacks, they still exhibit vulnerabilities at the implementation level, susceptible to hardware fault injections like the LeapFrog used in Rowhammer-based exploits.
In digital signature schemes, we find gadgets that produce several failure modes in the Open Quantum Safe Library. The most critical error is a bypass of the signature verification. Note, that while we experimented with Post-Quantum encryption schemes, theoretically LeapFrog gadgets should work on classical encryption schemes as well.
The “Magic Number Mismatch” column in Table IV highlights instances where the injected fault corrupts memory regions containing predefined magic numbers used for integrity checks. This mismatch signifies unintended memory corruption caused by the LeapFrog gadget, which can lead to unpredictable behavior or system crashes. According to Table IV, all signature schemes also contain gadgets for this failure mode, with Dilithium 3 containing the most number of LeapFrog gadgets.
Failures during key generation (“Key Gen. Fail”), signature generation (“Sig. Gen. Fail”), and signature verification (“Sig. Verif. Fail”) were also identified. Such failures can be exploited to disrupt normal cryptographic operations, resulting in denial-of-service (DoS) attacks or weakening cryptographic strength by producing invalid or insecure keys and signatures. All schemes contain this type of gadget.
The “Incorrect Verification” column denotes scenarios where invalid signatures are erroneously accepted as valid. This occurs when a LeapFrog gadget alters the control flow of the verification routine, enabling attackers to perform impersonation attacks by forging signatures that bypass standard validation checks.
Lastly, the “Verification Bypass” column in Table IV highlights instances where the signature verification routine can be entirely circumvented using LeapFrog gadgets. Similar to the TLS attack scenario described in Section 8.4, this allows an attacker to craft an invalid signature and have it accepted as valid at the client’s end. By flipping bits in the Program Counter (PC) values using the LeapFrog within the client’s memory space, the attacker effectively bypasses the signature verification routine. This vulnerability poses a significant security risk by enabling impersonation attacks and facilitating unauthorized access or actions within the system. Notably, Dilithium3 exhibits the highest number of LeapFrog gadgets for this threat, indicating a greater susceptibility to such attacks. An example of such a LeapFrog gadget in Dilithium is seen in Figure 11.
Scheme # Instructions Candidate Gadgets Magic Number Mismatch Key Gen. Fail Sig. Gen. Fail Sig. Verif. Fail Incorrect Verification Verification Bypass
Dilithium2 43474 17472 18 3 11 294 5 8
Dilithium3 42800 25010 36 6 22 654 10 20
Dilithium5 44171 18591 12 8 14 355 4 13
Falcon-512 60647 18641 18 16 50 106 4 16
Falcon-1024 60794 9360 8 8 22 40 2 7
Falcon-padded-512 60678 9396 8 9 21 41 2 6
Falcon-padded-1024 61128 9456 8 6 19 42 2 9
MAYO-1 43542 6924 8 3 4 33 2 8
MAYO-2 43533 6984 8 4 5 32 2 8
MAYO-3 46823 6996 8 3 4 35 2 6
MAYO-5 44302 6528 6 4 3 34 3 11
ML-DSA-44 43668 17616 0 3 1 21 1 5
ML-DSA-44-ipd 43338 8820 9 3 4 172 3 8
SPHINCS+-SHA2-128f 37109 8052 8 3 5 112 6 9
SPHINCS+-SHA2-128s 37379 8052 8 3 8 109 5 7
SPHINCS+-SHA2-192f 42143 8460 7 3 9 113 5 7
SPHINCS+-SHA2-192s 42535 8484 7 3 10 109 4 6
SPHINCS+-SHA2-256f 42437 8532 8 3 4 94 3 8
SPHINCS+-SHA2-256s 42758 8556 8 3 6 98 3 7
cross-rsdp-128-balanced 44024 10032 9 4 4 209 2 8
TABLE IV: Results from scans on the liboqs library, showing various issues encountered during signature operations for each digital signature scheme, along with the total number of assembly executions and candidate gadgets.
We find LeapFrog gadgets on FIPS 204 standard, also on other PQC digital signatures schemes, FALCON [20], MAYO [7], and CROSS [6].
Table IV summarizes LeapFrog gadgets found in the liboqs library for different PQC digital signature schemes. Compared to Dilithium, ML-DSA, the implementation of the FIPS 204 standard, had fewer LeapFrog gadgets, suggesting that its implementation might be more resilient to the specific fault attacks we conducted. However, this does not imply immunity, as the gadgets found were still capable of bypassing critical functions. The relatively lower number of vulnerabilities in ML-DSA could also be attributed to its simpler structure, which reduces the surface area for potential control flow subversion attacks.
We also evaluated SPHINCS+, a hash-based signature scheme standardized in FIPS 205. SPHINCS+ offers a different security foundation, relying on the hardness of hash-based constructions rather than lattice problems. While this scheme is robust against certain classes of attacks, our analysis uncovered several LeapFrog gadgets capable of bypassing signature verification. This suggests that even though the algorithm itself is designed to withstand quantum and classical cryptanalytic attacks, practical vulnerabilities arise due to implementation flaws that allow Rowhammer-based attacks to alter execution paths. Interestingly, the number of LeapFrog gadgets identified in SPHINCS+ varied significantly based on its parameter set, with some configurations being more resilient than others. This highlights the importance of parameter selection in mitigating the risk of physical attacks.
Figure 11: LeapFrog gadget detected in liboqs binary for Dilithium PQC Digital Signature Scheme. The PC value that fault is injected into, $addr_{src}$, is highlighted in blue. The new value after the fault injected, $addr_{dest}$, is highlighted in red. The fault is injected during the execution of the function call highlighted in green.
SPHINCS+ has more gadgets compared to the FALCON-1024 configuration, but in some configurations, it has fewer gadgets than FALCON-512, another selected algorithm that is not standardized. Overall, our findings indicate that there are generally more LeapFrog gadgets that enable bypassing signature verification compared to those that can falsely verify an invalid signature, indicating higher feasibility for DoS attacks with lower security impact compared to impersonation attacks.
@@ 8.4 TLS Handshake
Target Size #Inst.exec $d_{HD}$ Candidates
2 on 2 off
TLS 29KB 5328007 1 315 2493
2 2240 14413
3 21841 67421
TABLE V: Number of gadget candidates found in TLS scenario for fault models with different Hamming distances.
In a full end-to-end attack example, we illustrate the potency of the attack by applying it within a client/server authentication framework, specifically using OpenSSL for signature verification. Here, we consider a scenario where the attacker shares a physical computing space with the client. The goal of the attacker is to manipulate the client’s signature verification mechanism, causing it to erroneously validate a corrupted signature as genuine. This manipulation forms part of a broader man-in-the-middle strategy, aimed at deceiving the client into believing they are securely connected to the intended server.
@@
Figure 13: Probability distribution of bait page numbers.
We first profile the target memory to determine where the client places the PC value in recently deallocated pages. We use a process we call baiting [50] where we allocate a series of pages as the attacker. Then, we deallocate the flippy page, forcing the victim process to use recently-released flippy locations, where the PC values are stored. You can see our results in Figure 13 where we see that releasing 29 pages results in nearly a 30% chance of a PC value being stored in the next page released (which would be the flippy page).
We scan the client binary using MFS while the server is using the correct and incorrect private key. With step 2 on, we found 315 unique gadget candidates with $d_{H}=1$. When the server uses the correct key, the client binary terminates with return code 0, and when the server uses an incorrect key, the client returns with code 1. In step 5, we look for PC corruptions that cause the client to return with value 0, meaning it incorrectly authenticates the server. After the simulation, we found that one of the candidates was a LeapFrog gadget that caused false authentication of the malicious server with an incorrect key.
Then, we scan the client with step 2 off. With this mode, MFS detected 2493 gadget candidates with $d_{H}=1$. After the simulation steps, we verified that 21 of those candidates were LeapFrog gadgets that caused the client to return with 0, including the one found earlier. The number of candidates for different Hamming distance values is given in Table V.
The total time for the end-to-end attack to induce a successful misauthentication of the TLS handshake was 12 hours and 25 minutes, as seen in Table  VI. This time included profiling the system for the proper flippy pages with the correct offset, meaning the actual online time was around 2 hours. The experiment found a total of 1647 unique flippy pages, and over the course of the 2 hours of online attacking, we saw 2206 attacks where the program counter was baited into the correct page we were attacking before it flipped.
@@
TABLE VI: Results from the end-to-end attack on code using OpenSSL client/server signature verification with LeapFrog gadget
@@ 9 Countermeasures
Rowhammer Resistant Hardware
@@
Increasing the DRAM refresh rate is a commonly cited countermeasure to prevent Rowhammer attacks. Standard DRAM refresh is 64ms, meaning that a Rowhammer attack has 64ms to flip a bit before the row refreshes. Thus, a faster refresh rate will result in a shorter time window for the Rowhammer attack to be performed and should result in fewer flips. This is not an ideal solution, however, because a faster refresh rate will lead to worse power usage and performance overall. Alternative methods such as probabilistic row refresh [63] and parallel row refresh [66] are not available in consumer systems. Additionally, upgrading hardware to newer DDR5 technology has also proven to be ineffective [29].
@@
A novel countermeasure against Rowhammer attacks is the Randomized Row-Swap (RRS) method [52]. This approach fundamentally disrupts the spatial connection between aggressor and victim DRAM rows, thereby offering a robust defense against complex Rowhammer access patterns, including those not mitigated by victim-focused methods like the Half-Double attack. RRS operates by periodically swapping aggressor rows with randomly selected rows within the DRAM memory, limiting the potential damage to any single locality. While RRS can be implemented in conjunction with any tracking mechanism, its effectiveness has been demonstrated when paired with a Misra-Gries tracker, targeting a Row Hammer Threshold of 4.8K activations, akin to state-of-the-art attacks.
@@
Initial beliefs held that the Error Correcting Code (ECC) would serve as an effective defense against Rowhammer attacks. However, subsequent research has shown that ECC, despite its prevalence in server environments, falls short of a comprehensive solution. This inadequacy primarily arises due to ECC’s vulnerability to scenarios involving triple bit flips, a phenomenon well-documented in the literature [15]. Additionally, ECC, while standard in server-grade hardware, is typically absent in consumer-grade DRAM systems.
Adding nops To Code
@@
A mitigation against the LeapFrog attack specifically would be patching the source code or binary such that it is no longer vulnerable. Given the single-bit flip requirement of Rowhammer on the PC values, adding enough nops within the LeapFrog gadget to prevent instruction skips that only require a single-bit flip would potentially mitigate the attack. Adding nop instructions to source code is not trivial when the compiler optimizations are enabled since the compiler may reorder the critical parts in a different way, which makes the patch ineffective. A mitigation tool that adds nops to binary itself may overcome the compiler effect. Yet, adding new instructions to a binary will result in a change in the address of all the following instructions, which may introduce new LeapFrog gadgets. Therefore, the patched binary needs to be re-evaluated if the new version still has gadgets. Although a LeapFrog-aware compiler may potentially generate a LeapFrog proof binary, we claim it is not a sound and reliable approach.
Adding Redundancy to the Control Flow
Since LeapFrog gadgets are hard to mitigate in the source code and binary manually, we need a generic mitigation that can be implemented at the compiler level. The main target in LeapFrog is the program counter values that are temporarily pushed into the stack. Pushing multiple copies of the program counter to the stack and making sure the ultimate decision to return to an address is made on the combination of these copies would potentially make the attack impractical.
@@ 10 Conclusion
In this work, we introduced LeapFrog, a specific type of Rowhammer exploit that directly targets the control flow of programs by manipulating the Program Counter stored in the stack. This novel approach marks a significant shift in the understanding of Rowhammer threats, moving beyond traditional data integrity attacks to those that can alter program execution. Our successful demonstration of this attack in an OpenSSL TLS handshake scenario highlights its practical effectiveness and potential impact on widely used security protocols.
Furthermore, we proposed a systematic approach to identify LeapFrog gadgets in real-world software. Using our MFS analysis tool, we scanned multiple OpenSSL ciphers, Open Quantum Safe signature schemes, and machine learning classification algorithms and quantified the number of LeapFrog gadgets in this software. Even though the identification of vulnerable software is relatively straightforward thanks to our detection tool, mitigation of LeapFrog is not a trivial task since it is not transparent to the developers on a source code level. Instead, dedicated Rowhammer-resistant DRAM hardware or Rowhammer-aware compiler tools will be required to prevent LeapFrog attacks.